The base of the right prism is an equilateral triangle. The area of the base is 1,560 cm2. If the altitude of the prism is 5 meters,

____________________What is the base edge of the prism in cm?


____________________What is the lateral area of the prism in m2?


_____________________What is the volume of the prism in m3?
1
Expert's answer
2020-12-02T19:14:35-0500

Base area (B) = 1560 cm2 = 0.156 m2

Base area (B) "= \frac{\sqrt{3}}{4}a^2"

a - the base edge

"0.156 = \frac{\sqrt{3}}{4}a^2 \\

a^2 = \frac{0.624}{\sqrt{3}} \\

a = \sqrt{0.360} \\

a = 0.6 \; m = 60 \;cm"

The base edge of the prism is 60 cm.

Lateral area of the prism "= h \times P"

P - perimeter of triangle

In this case P

"P = 3a \\

P = 3 \times 0.6 = 1.8 \;m \\

h = 5 \;m"

Lateral area of the prism "= 5 \times 1.8 = 9 \; m^2"

Volume of prism

"V = B \times h \\

V = 0.156 \times 5 = 0.78 \;m^3"

The volume of the prism is 0.78 m3